40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 21 17:01:59 2022, Paco, Update, BHD, BH55 RFPD installed - part I IMG_3760.jpegIMG_3764.jpeg
    Reply  Thu Sep 22 19:51:58 2022, Anchal, Update, BHD, BH55 LSC Model Updates - part II 
       Reply  Fri Sep 23 19:04:12 2022, Anchal, Update, BHD, BH55 LSC Model Updates - part III 
          Reply  Thu Sep 29 18:01:14 2022, Anchal, Update, BHD, BH55 LSC Model Updates - part IV 
             Reply  Fri Sep 30 18:30:12 2022, Anchal, Update, ASS, Model Changes 
                Reply  Wed Oct 12 14:01:23 2022, Radhika, Update, ASS, Model Changes YARM_ASS.png
                   Reply  Wed Oct 12 18:18:16 2022, rana, Update, ASS, Model Changes 
    Reply  Fri Sep 23 14:10:19 2022, Radhika, Update, BHD, BH55 RFPD installed - part I IMG_3766.jpegIMG_3770.jpegIMG_3773.jpeg
       Reply  Fri Sep 23 18:31:46 2022, rana, Update, BHD, BH55 RFPD installed - part I 
          Reply  Mon Sep 26 11:39:37 2022, Paco, Update, BHD, BH55 RFPD installed - part II PXL_20220926_175010061.MOTION.jpgBH55_Transimpedance_Measurement.pdf
             Reply  Thu Oct 6 11:12:14 2022, Anchal, Update, BHD, BH55 RFPD installation complete 
Message ID: 17174     Entry time: Thu Oct 6 11:12:14 2022     In reply to: 17159
Author: Anchal 
Type: Update 
Category: BHD 
Subject: BH55 RFPD installation complete 

[Yuta, Paco, Anchal]

BH55 RFPD installation was still not complete until yesterday because of a peculiar issue. As soon as we would increase the whitening gain on this photodiode, we saw spikes coming in at around 10 Hz. Following events took place while debugging this issue:

  • We first thought that RFPD might be bad as we had just picked it up from what we call the graveyard table.
  • Paco fixed the bad connection issue at RF out and we confired RFPD transimpedance by testing it. See 40m/17159.
  • We tried changing the whitening filter board but that did not help.
  • We used BH55 RFPD to lock MICH by routing the demodulation board outputs to AS55 channels on WF2 board. We were able to lock MICH and increase whitening gain without the presence of any spikes. This ruled out any issue with RFPD.
  • Yuta and I tried swapping the whitening filter board but the problem persisted, which made us realize that the issue could be in the acromag that is writing the whitening gain for BH55 RFPD.
  • We combed through the /cvs/cds/caltech/target/c1iscaux/C1_ISC-AUX_LSCPDs.db file to check if the whitening gain DAC channels are written twice but that was not the case. But changing the scan rate of the whitening gain output channel did change the rate at which teh spikes were coming.
  • This proved that some other process is constantly writing zero on these outputs.
  • It tuned out that all unused channels of acromags for c1iscaux are still defined and made to write 0 through /cvs/cds/caltech/target/c1iscaux/C1_ISC-AUX_SPARE.db file. I don't think we need this spare file. If someone wants to use spare channels, they can quickly add it to dB file and restart the modbusIOC service on c1iscaux, it takes less than 2 minutes to do it. I vote to completely get rid of this file or atleast not use it in the cmd file.
  • After removing the violating channels, the problem with BH55 RFPD is resolved.

The installation of BH55 RFPD is complete now.

 

ELOG V3.1.3-